Suppose that when Keith’s income increases from RM28,000 to RM30,000 per year, his

purchases of good A increase from 4 units to 5 units.

i. Calculate the coefficient of income elasticity of demand.

ii. Interpret the value of the coefficient above.


1
Expert's answer
2022-02-18T09:02:16-0500

"IE=\frac{\frac{\Delta Q}{Q_1+Q_2}}{\frac{\Delta Y}{Y_1+Y_2}}=3.2>1."

Positive income elasticity of demand is associated with normal goods; an increase in income will lead to a rise in quantity demanded.


The elasticity of demand is greater than 1, it is a luxury good or a superior good.